George and Weezy received $31,700 of Social Security benefits this year ($12,300 for George; $19,400 for Weezy). They also received $5,450 of interest from jointly owned City of Ranburne Bonds and dividend income.

What amount of the Social Security benefits must George and Weezy include in their gross income under the following independent situations? (Do not round intermediate calculations. Round your final answers to the nearest whole dollar amount. Leave no answer blank. Enter zero if applicable.)

c. George and Weezy file married joint and receive $34,800 of dividend income from stocks owned by George.
Answer #1

1. If george george and Weezy file federal tax individually, then they have to pay tax as per following criteria:

Your adjusted gross income
+ Nontaxable interest
+ ½ of your Social Security benefits
= Your "combined income

Tax Liability for George-

Non Taxable Interest=(5450*123/317)= $2114.67

1/2 of Social Benefits= 12300/2= $6150

Combined Income= 2114.67+6150=$ 8264.67

If Combined income fall between $25000 and $34000 then tax will be 50% of benifits received.

So, tax in this situation will be 50% of $12300= $6150.

2.